# GLEE: A Unified Framework and Benchmark for Language-based Economic Environments

### v1

@@ 5 Conclusion
We present GLEE, a framework for evaluating the behavior of Large Language Models (LLMs) in language-based economic games. The goal of GLEE is to provide a comparative tool for assessing the performance of LLMs in various economic scenarios and enable their comparison to human players. We defined the game space within three main families of games: bargaining, negotiation, and persuasion, and introduced metrics to measure player performance. We developed a framework that allows for large-scale data collection from games between diverse LLMs and created an interface that facilitates the collection of data from games involving human players. Through this interface, we gathered data from 954K games between LLMs and from 3,405 games involving human players. The data is available for future research, which could advance the field of machine learning in language-based economic games, such as for predicting human decisions using artificial data and building more successful and human-like agents based on the metrics we define in GLEE.
Ethics Statement
Furthermore, our framework demonstrates the capability of collecting data from human players to understand the differences and similarities between LLMs and humans in economic environments. While this line of research has the potential for a strong scientific contribution, particularly in the field of behavioral economics, it also raises several ethical considerations. The collection of human data must be conducted with careful regulation and adherence to clear ethical guidelines, such as those established in venues like ICLR. The authors declare that there are no violations of the ICLR ethics code in our study. The entire process of data collection from human participants is elaborated in §3.2.
Acknowledgements
@@
Appendix A Game families through the lens of economic literature
@@
Bargining
As mentioned in §2, the standard bargaining model of Rubinstein (1982) consists of two players, Alice and Bob, engaging in alternating offers for a finite horizon ($T=\infty$) with commonly known discount factors $\delta_{A},\delta_{B}$. Our parameterization considers several additional degrees of freedom, such as finite vs. infinite time horizon, complete vs. incomplete information, and free language messages vs. structured and concise messages. In the standard model, Rubinstein (1982) showed that in the unique subgame-perfect equilibrium an agreement is reached in the first stage (i.e., utilities are not discounted), and the share of Alice is given by $Mp^{*}$, where $p^{*}=\frac{1-\delta_{B}}{1-\delta_{A}\delta_{B}}$. [^17] A subgame-perfect equilibrium is a strategy profile in which every player responds optimally in every hypothetical subgame of the game, including off-path scenarios that are not reached in practice. This solution concept can be seen as capturing a higher level of rationality compared to the alternative of Nash equilibrium. The case of finite horizon can be solved using backward induction, and typically results in a different outcome compared to the infinite case. As $T$ grows, the equilibrium outcome approaches the one of the infinite case. Extensions to incomplete information regarding the opponent’s discount factor are typically more challenging, and some of them are studied in the literature, e.g. Rubinstein (1985).
@@
If the seller is uncertain regarding the buyer’s valuation but has a prior belief distribution, a classical result by Harris & Raviv (1981) and Riley & Zeckhauser (1983) states that it is always optimal to sell the product at a fixed price. In contrast, if the seller does not have a prior belief over the buyer’s valuation, and instead aims to minimize regret, then an optional pricing policy will be to randomly choose a price from a carefully chosen distribution (Bergemann & Schlag, 2011).
@@
Our persuasion game follows the structure of a cheap talk game (Crawford & Sobel, 1982; Farrell & Rabin, 1996), where the sender (Alice) cannot commit to a signaling policy in advance, unlike in Bayesian persuasion models (Kamenica & Gentzkow, 2011). Under the particular payoff structure considered in our persuasion game, it is well-known that the cheap-talk game only admits a babbling equilibrium, i.e., an equilibrium in which all information is kept hidden (this is due to the strong misalignment of interests between the two players). In contrast, if the seller can commit to a signaling policy at the outset, as in Bayesian persuasion, then there exists a subgame-perfect equilibrium in which the seller commits to the following policy: When the product is of high quality, the seller recommends buying the product with probability $1$. When the product is of low quality, then the seller recommends with probability $q=\min\{\frac{p}{1-p}(v-1),1\}$. This policy is also incentive-compatible, in the sense that the buyer always buys the product when the seller recommends buying. [^18] In fact, the probability of lying $q$ is determined such that the buyer is indifferent upon receiving a recommendation, taking into account his belief updating, which relies on using Bayes’s law and knowing the seller’s committed policy. While the long-living buyer case is well-studied in the economic literature, such games often admit multiple equilibria, which makes the games difficult to analyze and predict (Kim, 1996; Aumann & Hart, 2003). As for the case of myopic buyers, Best & Quigley (2024) draws a connection between the repeated cheap talk game and the case of one-shot Bayesian persuasion, leading to an elegant analytical solution of the repeated game. Intuitively, the repetitive nature of the game induces a reputation effect, which plays a similar role to the commitment power in standard one-shot Bayesian persuasion.
Appendix B Data Collection System
@@
{”alice_gain”: The part that you will receive in your offer,
”bob_gain”: The part that Bob will receive in your offer,
”message”: The message you pass to Bob}
@@
“‘json {”alice_gain”: 900, ”bob_gain”: 100,
”message”: ”Let’s start fair. I’ll take the bigger share, but you get something too.”} “‘
@@
Answer with {”decision”: ”accept”} or {”decision”: ”reject”}
@@
“‘json {”decision”: ”reject”} “‘
@@
{”bob_gain”: The part that you will receive in your offer,
”alice_gain”: The part that Alice will receive in your offer,
”message”: The message you pass to Alice}
@@
“‘json {”bob_gain”: 500,
”alice_gain”: 500,
”message”: ”Let’s split it evenly. It’s the fairest way to start.”} “‘
@@
Answer with {”decision”: ”accept”} or {”decision”: ”reject”}
@@
“‘json {”decision”: ”accept”} “‘
@@
Appendix C Data Collection from Human Players
This appendix provides additional information on the human data collection interface described in §3.2, which facilitates data collection from GLEE games played between humans and LLMs.
C.1 Screenshots of the Data Collection Interface
@@
The structure of the application and the games consists of fixed parts and parts that vary between different game families. Each game starts with a screen where the player enters their name, followed by an instruction screen (Figure 2 in bargaining, Figure 6 in negotiation, and Figure 10 in persuasion). The instructions themselves can change depending on the type of game and the parameters of the game. On the instruction screen, there is a hidden prompt to enter a code word in the text box below, which is designed to filter out unfocused participants. If the player fails this test, they are taken to a screen that informs them of their failure. Otherwise, the game itself begins. In each round, both the human player and the LLM player perform an action, with the order depending on the game and configuration. An action could involve sending an offer to the other player (figure 3 in bargaining games, Figure 7 in negotiation games and Figure 11 in Persuasion games) or responding to the other player’s offer (for instance, Figure 4 in bargaining games, Figure 8 in negotiation games and Figure 12 in persuasion games). After both players have completed their actions, the human player is taken to a response screen (Figure 5), where he sees the decision of the LLM player. Afterward, if the game is still not over, the human player continues for another round. Once the game is finished, the player is taken to a quiz screen where they must answer a question related to the technical details of the game (Figure 9). If the human answer correctly, they are directed to the final screen (Figure 13), where they receive a code to enter on the mTurk website. If the player fails the final quiz, they do not receive a completion code and are taken to a screen that informs them of their failure in the quiz. In this case, we erase the game from our database.
C.1.1 Bargaining Games Screenshots
@@
C.1.2 Negotiation Games Screenshots
@@
C.1.3 Persuasion Games Screenshots
@@
C.2 Selection of Configurations for Human Data Collection
@@
C.3 Attention Checks for Human Players
@@
The first test appeared on the instruction screen. Toward the end of the instructions, a line requested players to write the code word ”sdkot” in a text box that appeared at the end of the instructions phase. Players who did not write this word were immediately disqualified and did not start the game, as they did not carefully read the instructions. A total of 412 players, representing 8.9% of those who began the game, failed this test.
@@
Appendix D Formulas used to train regression models
This appendix describes how the regression models used to analyze the data in §4 and Appendix E were built. The regression models are train to predict a specific metric based on the parameters that define the configuration. To prevent the regression models from learning relationships directly from the configuration settings themselves, all parameters are treated as categorical variables.
The features included in the regression models are classify according to different levels of complexity, with each level describing the number of interactions between the parameters defining the feature.
The following code demonstrates the creation of the set of the features. The ”:” symbol represents an interaction between the parameters used to create the features.
1features += [p for p in game_parameters]
2features += ["alice_model", "bob_model"]
3features += ["alice_model:bob_model"]
4features += [f"{p}:alice_model" for p in game_parameters]
5features += [f"{p}:bob_model" for p in game_parameters]
6features += [f"({p1}:{p2})" for p1 in game_parameters for p2 in game_parameters if p1 != p2]
7features += [f"({p1}:{p2}):alice_model" for p1 in game_parameters for p2 in game_parameters if p1 != p2]
8features += [f"({p1}:{p2}):bob_model" for p1 in game_parameters for p2 in game_parameters if p1 != p2]
9features = "+".join(features)
10formula = f"{metric}~{features}"
After constructing the feature matrix, identical features may emerge212121for example, in a case where Language Model A, playing as Alice, only played games with Language Model B, playing as Bob. In such a case, the feature ”Is M=100 and Alice played by Language Model A” and the feature ”Is M=100, Alice played by Language Model A, and Bob played by Language Model B” will always be identical.. We are remove the more complex features from the model to allow us to measure the true contribution of each individual feature. The resulting formula is used to train the regression models.
Appendix E Graphs
This appendix presents graphs depicting the effect of each parameter on player gains, fairness, and efficiency in the games, depending on the language model that played the games. The graphs include 95% confidence intervals. [^22] The confidence intervals presented for human behavior are significantly larger compared to those for the language models, as the amount of data collected from humans is smaller by three orders of magnitude. The impact of each parameter is shown in a separate figure, which contains sub-graphs illustrating the effect of changing the parameter’s value on self-gain, efficiency, and fairness for each language model, both for Alice and Bob. Table 6 serves as a key for quick navigation to each of the figures.

### v2

@@ 5 Conclusion
We present GLEE, a framework for evaluating the behavior of Large Language Models (LLMs) in language-based economic games. The goal of GLEE is to provide a comparative tool for assessing the performance of LLMs in various economic scenarios and enable their comparison to human players. We defined the game space within three main families of games: bargaining, negotiation, and persuasion, and introduced metrics to measure player performance. We developed a framework that allows for large-scale data collection from games between diverse LLMs and created an interface that facilitates the collection of data from games involving human players and LLMs.
Through this interface, we gathered data from a variety of economic games, including both LLM vs. LLM and human vs. LLM interactions. Our analysis uncovered several notable phenomena: the varying impact of market parameters on economic outcomes, the interdependence between the game outcome and the type of LLMs used for both players, and the deviations of human performance from that of LLMs across different game families. We believe that GLEE offers a valuable foundation for advancing interdisciplinary research into the economic implications of LLMs, enabling systematic exploration of how AI agents and human decision-makers strategically interact and influence economic outcomes.
Acknowledgments and Disclosure of Funding
@@
Appendix A How GLEE Differs from Prior Benchmarks
To put in context the contribution of GLEE within the broader research landscape, we compare it to several recent benchmarks designed to evaluate the performance of LLMs in strategic settings. While these benchmarks differ in their choice of games, evaluation metrics, and interaction modalities, none focus on multi-turn economic games involving natural language communication. GLEE addresses this gap by offering a unified, richly parameterized framework for studying LLMs in sequential, language-based economic interactions. The following comparisons highlight where GLEE introduces new capabilities and how it complements or extends existing benchmarks.
LLMs as Rational Players [15]. Test the rationality of LLM in the light of game theory, using three simple one-shot games with no interaction between the players. GLEE supplies multi-round negotiation/bargaining/persuasion with free-text chat and direct human–agent comparison.
GAMA-Bench [23]. Introduce 8 single-turn games studied in game theory to test LLM rationality. The games are not economic games with little interaction between players. GLEE targets language–centric games, with games extensively studied in economic literature and extensive human trials.
GTBench [14]. Introduces a benchmark to test LLM rationality using board and card games. These games are not economic games, and natural language interaction does not happen. GLEE emphasis is on natural language communication (e.g., concessions, persuasion) within purely economic games.
Game theoric LLM [22]. Measures LLM behaviour on different multi-turn non-economic games extensively studied in game theory. The games they use have no extensive parameterization, and therefore few variants. GLEE is based on three economic games with heavy parameterization, allowing testing LLMs on hundreds of variants of the same game.
TMGBench [48]. Introduces an exhaustive benchmark with 144 $2\times 2$ normal-form games, but these games are not economic games, are often synthetic, single-shot, and without communication between the players. While GLEE covers fewer games, it captures real dynamics with a multi-turn and natural language communication between the player.
CompeteAI [51]. Introduce a one-competition world with multiple players; the metrics used are qualitative (LLM judgements). GLEE generalises across three economic game families with multiple parametrizations, using quantitative metrics derived from game theory research.
We summarize key differences in Table 4, highlighting where GLEE introduces new capabilities.
Table 4: Comparison of benchmarks evaluating LLMs in strategic games. GLEE stands out by focusing on multi-turn, natural language, economic interactions with extensive human evaluation and openness.
Framework Multi turns NL Communication Human Evaluation
GLEE (ours)
Rational Players [15]
GAMA-Bench [23]
GTBench [14]
Game theoric LLM [22] Structured
TMGBench [48]
CompeteAI [51] Indirect
Appendix B Sequential vs. Simultaneous Games
A key feature of all these games is that they are sequential, meaning players take turns acting rather than acting simultaneously. This makes communication, and in particular language-based communication, a crucial element of the interaction. In sequential games, communication is typically direct, with players able to attach messages to their actions (e.g., in the bargaining example, the textual message is attached to the proposal). In contrast, in simultaneous games, communication tends to be indirect, as players act independently and typically learn to cooperate by observing and reacting to past actions rather than through direct message exchange. A prime example of a simultaneous, language-based economic interaction is the competition among web publishers in search engines. Publishers simultaneously create content (usually in the form of textual documents) for their websites, and then gradually learn and adapt based on outcomes, such as their relative search engine rankings and exposure. However, they usually lack the opportunity for real-time communication or coordination during this process. These simultaneous, language-based interactions are well-studied within the information retrieval community, and the induced publishers’ game is indeed modeled as a simultaneous game (see e.g. [37, 6, 26, 21, 29, 30, 33], in which SEO competitions are modeled as simultaneous games). Our work is complementary, focusing on sequential settings where direct, language-based communication is prevalent.
Appendix C Game Families Through the Lens of Economic Literature
@@
Bargaining
As mentioned in §2, the standard bargaining model of [40] consists of two players, Alice and Bob, engaging in alternating offers for a finite horizon ($T=\infty$) with commonly known discount factors $\delta_{A},\delta_{B}$. Our parameterization considers several additional degrees of freedom, such as finite vs. infinite time horizon, complete vs. incomplete information, and free language messages vs. structured and concise messages. In the standard model, Rubinstein [40] showed that in the unique subgame-perfect equilibrium an agreement is reached in the first stage (i.e., utilities are not discounted), and the share of Alice is given by $Mp^{*}$, where $p^{*}=\frac{1-\delta_{B}}{1-\delta_{A}\delta_{B}}$. [^14] A subgame-perfect equilibrium is a strategy profile in which every player responds optimally in every hypothetical subgame of the game, including off-path scenarios that are not reached in practice. This solution concept can be seen as capturing a higher level of rationality compared to the alternative of Nash equilibrium. The case of a finite horizon can be solved using backward induction, and typically results in a different outcome compared to the infinite case. As $T$ grows, the equilibrium outcome approaches the one of the infinite case. Extensions to incomplete information regarding the opponent’s discount factor are typically more challenging, and some of them are studied in the literature, e.g. [41].
@@
If the seller is uncertain regarding the buyer’s valuation but has a prior belief distribution, a classical result by Harris and Raviv [19] and Riley and Zeckhauser [39] states that it is always optimal to sell the product at a fixed price. In contrast, if the seller does not have a prior belief over the buyer’s valuation, and instead aims to minimize regret, then an optional pricing policy will be to randomly choose a price from a carefully chosen distribution (7).
@@
Our persuasion game follows the structure of a cheap talk game [12, 16], where the sender (Alice) cannot commit to a signaling policy in advance, unlike in Bayesian persuasion models [24]. Under the particular payoff structure considered in our persuasion game, it is well-known that the cheap-talk game only admits a babbling equilibrium, i.e., an equilibrium in which all information is kept hidden (this is due to the strong misalignment of interests between the two players). In contrast, if the seller can commit to a signaling policy at the outset, as in Bayesian persuasion, then there exists a subgame-perfect equilibrium in which the seller commits to the following policy: When the product is of high quality, the seller recommends buying the product with probability $1$. When the product is of low quality, then the seller recommends with probability $q=\min\{\frac{p}{1-p}(v-1),1\}$. This policy is also incentive-compatible, in the sense that the buyer always buys the product when the seller recommends buying. [^15] In fact, the probability of lying $q$ is determined such that the buyer is indifferent upon receiving a recommendation, taking into account his belief updating, which relies on using Bayes’s law and knowing the seller’s committed policy. While the long-living buyer case is well-studied in the economic literature, such games often admit multiple equilibria, which makes the games difficult to analyze and predict [25, 5]. As for the case of myopic buyers, [8] draws a connection between the repeated cheap talk game and the case of one-shot Bayesian persuasion, leading to an elegant analytical solution of the repeated game. Intuitively, the repetitive nature of the game induces a reputation effect, which plays a similar role to the commitment power in standard one-shot Bayesian persuasion.
Appendix D LLM Data Collection
D.1 Game Management System Interface
In this appendix, we present the main features of our game management system through a series of screenshots. The system facilitates data collection and analysis from the games described in §2, offering a user-friendly and customizable interface.
The main interface of the system (Figure 2) provides three primary options: starting a new data collection, resuming a previously interrupted collection, and analyzing results from past experiments. When initiating a new data collection, users first select the game family and the participating LLMs (Figure 3), followed by defining the set of configurations to be executed (Figure 4).
Once the data collection is complete, users can move to the analysis phase. The data analysis module starts with selecting the data to be analyzed (Figure 5). After selecting the relevant data, users can use the parameter impact viewer (Figure 6) to visualize how game parameters influence various metrics (see §3). Additionally, the system provides a statistics table (Figure 7) that summarizes key characteristics of the collected data.
The system does not require any specialized hardware: data collection and analysis can be performed efficiently using a single CPU.
Figure 2: Main interface: Start a new data collection, resume a previous one, or analyze results from past experiments.
Figure 3: New data collection: Selecting game family and LLMs.
Figure 4: New data collection: Setting up configurations.
Figure 5: Data analysis: Selecting datasets for analysis.
Figure 6: Data analysis: Visualizing the impact of game parameters on various metrics.
Figure 7: Data analysis: Statistics table summarizing key characteristics of the collected data.
D.2 Conversation Structures
@@
{"alice_gain": The part that you will receive in your offer,
"bob_gain": The part that Bob will receive in your offer,
"message": The message you pass to Bob}
@@
“‘json {"alice_gain": 900, "bob_gain": 100,
"message": "Let’s start fair. I’ll take the bigger share, but you get something too."} “‘
@@
Answer with {"decision": "accept"} or {"decision": "reject"}
@@
“‘json {"decision": "reject"} “‘
@@
{"bob_gain": The part that you will receive in your offer,
"alice_gain": The part that Alice will receive in your offer,
"message": The message you pass to Alice}
@@
“‘json {"bob_gain": 500,
"alice_gain": 500,
"message": "Let’s split it evenly. It’s the fairest way to start."} “‘
@@
Answer with {"decision": "accept"} or {"decision": "reject"}
@@
“‘json {"decision": "accept"} “‘
@@
D.3 LLM List
Table 5 presents the 13 LLMs used for data collection (see 3), organized by model name, developer, and release year.
Table 5: Large Language Models used for data collection.
Developer Model Name Release Year
Anthropic Claude 3.5 Sonnet 2024
Claude 3.7 Sonnet 2024
Google DeepMind Gemini 1.5 Flash 2024
Gemini 1.5 Pro 2024
Gemini 2.0 Flash 2025
Gemini 2.0 Flash Lite 2025
OpenAI GPT-4o 2024
GPT-4o Mini 2024
GPT-O3 Mini 2024
xAI Grok 2 2024
Meta LLaMA 3.1–40.5B 2024
LLaMA 3.3–70B 2024
Mistral AI Mistral Large 2024
D.4 Statistics
Table 6 contains statistics on the data we collected, presented in §3, which constitutes a contribution of the paper.
Table 6: Statistics of data collected by family.
Game Type Games Decisions Messages Words
Bargaining 33.7K 92.8K 45.8K 1.14M
Persuasion 13.5K 402K 140K 6.95M
Negotiation 32.9K 92.3K 44.4K 1.69M
Total 80.1K 587K 230K 9.77M
Appendix E Human Data Collection
This appendix provides information on the human data collection interface described in §3, which facilitates data collection from GLEE games played between humans and LLMs.
One of the main objectives of collecting data from language games is to compare the behavior of LLMs with human behavior in economic and strategic situations. To facilitate this comparison, we developed an interface that allows human players to play all the language games that can be defined using GLEE. The interface transforms the various prompts presented to LLMs into user-friendly screens for human participants, displaying the prompt and requesting them to send messages and make decisions. [^16] Since human players are accustomed to being addressed by their first name (rather than as Alice or Bob), we asked them to enter their name at the beginning of the game and referred to them by their name throughout the game. The player’s name was the only difference between the human player and the LLM-based player. Through this interface, we enable human players to take on the role of one of the players while the other player is a pre-selected LLM. The interface is one of the contributions of this paper. Screenshots of the interface can be found in Appendix E.1.
The interface, developed using oTree (Chen et al. 10), enables integration with Amazon’s crowdsourcing platform, mTurk, [^17] https://www.mturk.com/. through which we recruited 3,405 players who participated in various configurations against Gemini-1.5-flash. We chose Gemini-1.5-flash since it demonstrated strong performance compared to other LLMs and allowed comprehensive data collection due to its low usage cost. Since we aimed to collect multiple games from each configuration, we had to select a limited set of configurations for human participants to play. The process of selecting these sets is described in Appendix E.2. We collected human data from 195 different configurations: 78 of which were bargaining games, 60 of which were negotiation games, and 57 were persuasion games.
Each human player was allowed to play one game every 12 hours from each family of games. Human players were paid a base rate calculated at \$6 per hour, plus an average bonus of \$6 per hour. In total, we paid \$2,245 to all players for their participation in the games. Bonuses were dependent on the configuration the human players played and their success in the game. The average bonus was known to players at the start of the game, aiming to encourage serious gameplay. To ensure players remained focused and made thoughtful decisions, we conducted two attention checks during the experiment, detailed in Appendix E.3. Players who failed in one of the attention checks were excluded from the dataset. To reflect the real-world significance of the magnitude of product prices (the parameter M) in each family of games, we defined the bonus as dependent on M: in configurations where M = $10^{2}$, the average bonus was \$3 per hour; where M = $10^{4}$, the average bonus was \$6 per hour; and where M = $10^{6}$, the average bonus was \$9 per hour.
E.1 Screenshots of the Data Collection Interface
@@
The structure of the application and the games consists of fixed parts and parts that vary between different game families. Each game starts with a screen where the player enters their name, followed by an instruction screen (Figure 8 in bargaining, Figure 12 in negotiation, and Figure 16 in persuasion). The instructions themselves can change depending on the type of game and the parameters of the game. In the middle of the game’s rules on the instruction screen, there is a hidden prompt directing participants to enter a code word in the text box below. This test is designed to filter out unfocused participants. If the player fails this test, they are directed to a screen informing them of their failure and will not participate in the game. Otherwise, the game begins. In each round, both the human player and the LLM player perform an action, with the order depending on the game and configuration. An action could involve sending an offer to the other player (Figure 9 in bargaining games, Figure 13 in negotiation games and Figure 17 in Persuasion games) or responding to the other player’s offer (for instance, Figure 10 in bargaining games, Figure 14 in negotiation games and Figure 18 in persuasion games). After both players have completed their actions, the human player is taken to a response screen (Figure 11), where he sees the decision of the LLM player. Afterward, if the game is still not over, the human player continues for another round. Once the game is finished, the player is taken to a quiz screen where they must answer a question related to the technical details of the game (Figure 15). If the human answer correctly, they are directed to the final screen (Figure 19), where they receive a code to enter on the mTurk website. If the player fails the final quiz, they do not receive a completion code and are taken to a screen that informs them of their failure in the quiz. In this case, we erase the game from our database.
E.1.1 Bargaining Games Screenshots
@@
E.1.2 Negotiation Games Screenshots
@@
E.1.3 Persuasion Games Screenshots
@@
E.2 Selection of Configurations for Human Data Collection
@@
E.3 Attention Checks for Human Players
@@
The first test appeared on the instruction screen. Toward the end of the instructions, a line requested players to write the code word "sdkot" in a text box that appeared at the end of the instructions phase. Players who did not write this word were immediately disqualified and did not start the game, as they did not carefully read the instructions. A total of 412 players, representing 8.9% of those who began the game, failed this test.
@@
Appendix F Additional Details on the Statistical Analysis
F.1 Input of Linear Regression
To support the statistical analysis presented in §3, we employed linear regression models. Each model was trained to predict a single outcome metric (fairness, efficiency, Alice’s self-gain, or Bob’s self-gain) within a specific family. The input features included all game-defining parameters (as listed in Table 1), along with identifiers for the two players involved in the game.
To represent player identity, we introduced two additional features, each indicating one of the players. [^20] When analyzing model interactions (e.g., in Tables 21, 22, 23), we replaced these two features with a single feature representing the pairwise identity of the two LLMs. Within each game family, we combined the parameters defining the market (T, CI and MA/MT - see Table 1) into a single parameter named market, representing the interaction between the three original parameters. We do so because their interaction defines the structure of the market itself, and the total number of combinations was small enough to allow reliable modeling. Thus, for example, in bargaining games, the market parameter became a vector with eight entries, corresponding to each possible combination arising from the Cartesian product of these three market-defining parameters.
Each parameter was represented using a one-hot encoding vector, where each possible value of the parameter was assigned a distinct entry. The entry corresponding to the actual value of the parameter was set to 1, while all other entries were set to 0. For instance, the parameter $M$ was split into three distinct parameters: $M=10^{2}$, $M=10^{4}$ and $M=10^{6}$, each taking the value of 1 if the respective condition held and 0 otherwise. Figure 20 illustrates this feature representation.
Figure 20: Illustration of feature encoding for linear regression in a bargaining game. Each row corresponds to a single bargaining game instance, encoded as a binary feature vector. The input includes game-defining parameters ($\delta_{A}$, $\delta_{B}$, $M$), alongside a composite market feature derived from the interaction of three specific parameters: $T$, $CI$, and $MA$. This market feature is one-hot encoded with eight possible values, representing all combinations of the three. Player identities (Alice and Bob) are encoded using separate one-hot vectors with 13 possible values each. The binary vectors at the bottom of the figure illustrate how all components are concatenated into a single input row, where each 1 indicates the active value of a categorical parameter for this particular bargaining game instance.
F.2 Baseline Models
Table 8 presents the default parameter values used in our statistical analysis (see §3). These values serve as the reference points for estimating the impact of each parameter value on the target metrics. The selection of default values is based on the simplest model in game theory.
Table 8: Default parameter values used in the statistical analysis. $T=\infty$ means a very large value of $T$, unknown to the players. CI = Complete Information. MA = Textual messages allowed.